OpenStudy (anonymous):

Please Help..... Find the value of the annuity.. A1=$1000, I=0.05, N=25

OpenStudy (anonymous):

\[a _{1}=1000, I=0.05, n=25\]

OpenStudy (campbell_st):

ok.. so the future value formula is \[A = \frac{A_{1}[(1 + I)^n -1]}{I}\] so in your question A1 = 1000 I = 0.05 and n = 25 substitute then and evaluate.

OpenStudy (anonymous):

\[1000=\frac{ 1000\left[ \left( 1+0.05 \right)^{25}-1 \right] }{ 0.05 }\] How do i solve it? @campbell_st

OpenStudy (campbell_st):

just type it into a calcuator as its written.

OpenStudy (anonymous):

how do i raise it to the 25th power? @campbell_st

OpenStudy (campbell_st):

your calculator should be able to handle it...

OpenStudy (campbell_st):

do you have a scientific of graphing calculator..?

OpenStudy (anonymous):

no i do not @campbell_st

OpenStudy (campbell_st):

ok... I got $47727.10 here is an online calculator so you can check it http://web2.0calc.com/